Caspion Corporation makes and sells a product called a Miniwarp. One Miniwarp requires 2.5 kilograms of the raw material JuriCaspion Corporation makes and sells a product called a Miniwarp. One Miniwarp requires 2.5 kilograms of the raw material Jurislon. Budgeted production of Miniwarps for the next five months is as follows: August 23,500 units September 22,200 units October 23,600 units November 24,800 units December 24,500 units The company wants to maintain monthly ending inventories of Jurislon equal to 25% of the following month's production needs. On July 31, this requirement was not met since only 11,700 kilograms of Jurislon were on hand. The cost of Jurislon is $27 per kilogram. The company wants to prepare a Direct Materials Purchase Budget for the next five months. The total cost of Jurislon to be purchased in August is:
Answer #1

Purchase quantity = raw material needed for production+desired ending inventory-Beginning inventory

= (23500*2.50)+(22200*2.5*25%)-11700

Purchase quantity = 60925

Purchase cost = 60925*27 = $1644975

So answer is c) $1644975